4a. If you have two assets with expected returns 10% and 5%, respectively, what is the percentage you have to invest in every assets in order to get an expected return of 8%? What would be the risk on that portfolio if the covariance between the two assets is zero? Standard Deviation is 56.66%
Answer #1

Weight of asset A = x
Weight of asset B = 1-x

Expected Return = x*10%+(1-x)*5% = 8%
5%x = 3%
x = 60%
1-x = 40%
Percentage of Asset 1 = 60% and Percentage of Asset 2 = 40%

Standard deviation is not given Hence overall standard deviation cannot be calculated

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well